# Counterfactual Fairness Evaluation of Machine Learning Models on Educational Datasets

### v1

@@ 2.3 Algorithmic Fairness in Education
Table 1: Feature descriptions of Law School and OULAD datasets. Student Performance dataset descriptions are provided in Table 6 of Appendix 0.A.
@@ 3.1 Educational Datasets
We use publicly available benchmark educational datasets for fairness presented in [26], which introduces four educational benchmark datasets for algorithmic fairness. Datasets are Law School111github.com/mkusner/counterfactual-fairness[46], Open University Learning Analytics Dataset (OULAD) [^2] https://archive.ics.uci.edu/dataset/349/open+university+learning+analytics+dataset[24] and Student Performance in Mathematics and Portuguese language [^3] https://archive.ics.uci.edu/dataset/320/student+performance[8]. Selected features and its descriptions of datasets are presented in Table 1 and the summary of tasks and selection of sensitive attributes are outlined in Table 2.
@@
The Student Performance dataset describes students’ achievements in Mathematics and Portuguese language subjects in two Portuguese secondary schools during 2005-2006. The dataset provides details about students’ demographics, and family backgrounds such as parent’s jobs and education level, study habits, extracurricular activities, and lifestyle. We select gender as the sensitive attribute and G3 as the target for the regression problem. Feature description of the dataset is presented in Appendix 0.A.
@@
Figure 2: Partial DAGs of the estimated causal model for educational datasets, showing only the sensitive attribute, its descendants, and the target variable. See Appendix A for full graphs. Each sub-graph is not used for implementing counterfactually fair models; only the remaining features are included.
@@ 3.2 Structural Causal Model of Educational Dataset
To construct the SCM of OULAD and the Student Performance dataset, we use a causal discovery algorithm, Linear Non-Gaussian Acyclic Model (LiNGAM). The algorithm estimates a causal structure of the observational data of continuous values under linear-non-Gaussian assumption[41]. From the estimated causal model, we filtered DAG weights that are under the 0.1 threshold.
@@ 3.4 Experiment Details
We evaluate the counterfactual fairness of machine learning models on both regression and classification models. We selected the four most utilized machine learning models in the algorithmic fairness literature[19]. We choose Linear Regression(Logistic Regression for classification), Multilayer Perceptron(MLP), Random Forest(RF), and XGBoost(XGB)[5]. For KDE plot visualizations, we used a linear regression model for regression and MLP for classification.
@@ 5 Discussion
The causal model of Student Performance is presented in Fig. 2(c) and 2(d). The estimated causal model shows potential gender-based influences in study habits, social behaviors, and alcohol consumption to academic performance. Foremost, gender have an indirect causal relationship on G3. For both datasets, gender directly influences studytime, and studytime directly influences G1. For Mathematics, gender directly impacts studytime, freetime, goout and Dalc, but not goout for Portuguese. Differences in goout and alcohol consumption(Dalc and Walc) show that the factors influencing student performance differ between Math and Portuguese, demonstrating the importance of considering subject-specific causal models in education.
@@ 6.0.1 Acknowledgements
Appendix 0.A Appendix
Construction of causal structural model(SCM) is crucial for implementing counterfactual fairness. Thus, we provide a estimated SCM inferred from each dataset through LiNGAM algorithm. We filtered out edges with absolute weights lower than 0.1. These causal models are used for sampling counterfactual instances. For fitting a counterfactually fair model, we excluded direct and indirect descendants of the sensitive feature for each dataset.

### v2

@@ 2.3 Algorithmic Fairness in Education
Table 1: Feature descriptions of Law School and OULAD datasets. Student Performance dataset descriptions are provided in Table 6 of Appendix A.
@@ 3.1 Educational Datasets
We use publicly available benchmark educational datasets for fairness presented in [26], which introduces four educational benchmark datasets for algorithmic fairness. Datasets are Law School111github.com/mkusner/counterfactual-fairness[46], Open University Learning Analytics Dataset (OULAD) [^2] https://archive.ics.uci.edu/dataset/349/open+university+learning+analytics+dataset[24] and Student Performance in Mathematics and Portuguese language [^3] https://archive.ics.uci.edu/dataset/320/student+performance[8]. Refer to Table 1 and Table 6 for description of dataset features used in the experiment. The summary of tasks and selection of sensitive attributes are outlined in Table 2.
@@
The Student Performance dataset describes students’ achievements in Mathematics and Portuguese language subjects in two Portuguese secondary schools during 2005-2006. The dataset provides details about students’ demographics, and family backgrounds such as parent’s jobs and education level, study habits, extracurricular activities, and lifestyle. We select gender as the sensitive attribute and G3 as the target for the regression problem. Feature description of the dataset is presented in Appendix A.
@@
Figure 2: Partial DAGs of the estimated causal model for educational datasets, showing only the sensitive attribute, its descendants, and the target variable. See Appendix B for full graphs. Each sub-graph is not used for implementing counterfactually fair models; only the remaining features are included.
@@ 3.2 Structural Causal Model of Educational Dataset
To construct the SCM of OULAD and the Student Performance dataset, we use a causal discovery algorithm, Linear Non-Gaussian Acyclic Model (LiNGAM) [41]. The algorithm estimates a causal structure of the observational data of continuous values under linear-non-Gaussian assumption. From the estimated causal model, we filtered DAG weights that are under the 0.1 threshold.
@@ 3.4 Experiment Details
We evaluate the counterfactual fairness of machine learning models on both regression and classification models. We selected the four most utilized machine learning models in the algorithmic fairness literature[19]. We choose Linear Regression(LR; Logistic Regression for classification), Multilayer Perceptron(MLP), Random Forest(RF), and XGBoost(XGB)[5]. For KDE plot visualizations, we used a linear regression model for regression and MLP for classification.
@@ 5 Discussion
The causal model of Student Performance is presented in Fig. 2(c) and Fig. 2(d). The estimated causal model shows potential gender-based influences in study habits, social behaviors, and alcohol consumption to academic performance. Foremost, gender have an indirect causal relationship on G3. For both datasets, gender directly influences studytime, and studytime directly influences G1. For Mathematics, gender directly impacts studytime, freetime, goout and Dalc, but not goout for Portuguese. Differences in goout and alcohol consumption(Dalc and Walc) show that the factors influencing student performance differ between Math and Portuguese, demonstrating the importance of considering subject-specific causal models in education.
@@ 6.0.1 Acknowledgements
Appendix A Feature Description of Student Performance Dataset
@@
Appendix B Complete SCMs of Datasets
Construction of causal structural model(SCM) is crucial for implementing counterfactual fairness. Thus, we provide a estimated SCM inferred from each dataset through LiNGAM algorithm[41]. We filtered out edges with absolute weights lower than 0.1. These causal models are used for sampling counterfactual instances. For fitting a counterfactually fair model, we excluded direct and indirect descendants of the sensitive feature for each dataset.